# Generative Models in Decision Making: A Survey

### v2

@@ 4.2 Generative Models as Modeler
Unlike previous method S4RL [195] or RAD [196], which augments the states or the input spaces with slight perturbation to ensure the consistency, generative models take a different approach. Instead of merely modifying existing inputs, generative models focus on learning the underlying patterns within a dataset, thereby generating new data that closely resembles the original distribution [197].
In other words, a generative model processes a given set of input data and analyzes it to uncover the inherent patterns and structures of the data points [198]. In this context, generative models can be viewed as powerful ”modelers”, capable of capturing complex relationships within data and creating new and similar data that reflect these intrinsic relationships. This means that they can be employed to extract data representations for downstream tasks, such as classification or regression.
For example, generative models are capable of generating new images that resemble the images in the training dataset [199]. The model analyzes the features and patterns of the original images and uses this information to produce new images that mirror the style and content of the originals, just like DALL·E [1] does. Similarly, generative models can be used to create new text similar in style and content to the training corpus. Beyond image and text generation, they are useful for creating synthetic data for purposes like data augmentation, privacy protection, and simulations.
Overall, a generative model can be viewed as a ”modeler” because it learns the underlying patterns from a dataset [200] and generates new data that closely resembles the original data.
@@ 4.2.1 EBMs as Modeler
As ”models”, EBMs model data through an energy function that captures the underlying structure. Unlike other generative models, EBMs do not rely on explicit probabilities distributions but instead use an energy landscape to describe the data. Once trained, EBMs can generate new data by minimizing the energy, ensuring that the generated samples align with the learned patterns of the original dataset.
Boltzmann Machines (BM)[33] uses contrastive divergence techniques to generate samples that match the training data’s distribution. Similarly, DEBMs [121] extend this approach by using energy-based models to learn a distribution over state-action pairs, enabling reinforcement learning agents to generate and select optimal actions through contrastive divergence. Score-Based Generative Models (SGMs) [100] use score matching techniques to learn the gradients of the data distribution, enabling the generation of samples by refining noisy data towards the target distribution.
@@ 4.2.2 GANs as Modeler
GANs serve as models by learning the underlying patterns through their unique adversarial framework, which consists of two parts: the generator and the discriminator. Synthetic data is generated by the generator, with the discriminator assessing its realism, which in turn drives the generator to refine its output.
EGAN [134] serves as a modeler by leveraging the relationship between states and actions to learn the underlying patterns of the environment, which enables the pre-training of the agent and accelerates the learning process. Similarly, State2Pixel (S2P) [135] synthesizes raw pixel images from the agent’s state, effectively bridging the gap between the state space and the image domain in reinforcement learning (RL) algorithms. Additionally, S2P facilitates virtual exploration of the latent image distribution through model-based transitions in the state space.
@@ 4.2.3 VAEs as Modeler
VAEs are well-suited as modelers due to their capacity in learning a probabilistic mapping from the data space to the latent space, capturing the underlying variations in the data.
Han & Kim [145] represent the subspaces of the dataset by using the variational autoencoder. Subsequently, the VAE decoder is used to extract new data from the latent space and maps it back to the original one.
@@ 4.2.4 Normalizing Flows as Modeler
Normalizing Flows (NF) serve as powerful tools for modeling complex, high-dimensional data distributions by applying a series of invertible transformations. This flexibility enables NF to capture intricate data structures, making them particularly useful for tasks such as density estimation, variational inference, and posterior distribution modeling.
For instance, Non-linear Independent Component Estimation (NICE) [90] applied NF to model complex distributions by applying a series of invertible transformations, allowing it to represent high-dimensional data distributions flexibly. Rezende and Mohamed [83] further expanded the use of NF by applying them in variational inference to model complex posterior distributions. By applying NF to parameterize the posterior, they were able to construct flexible, tractable distributions that could be used for efficient Bayesian inference. This method helped overcome challenges associated with approximating posterior distributions, offering a scalable approach to variational inference in high-dimensional spaces. Müller et al. [201] enhanced NF’s flexibility by introducing piecewise-polynomial coupling transformations, which enhanced NF’s ability to model more complex distributions. By applying NF to parameterize the posterior, they were able to construct flexible, tractable distributions that could be used for efficient Bayesian inference. This method helped overcome challenges associated with approximating posterior distributions, offering a scalable approach to variational inference in high-dimensional spaces.
@@ 4.2.5 Diffusion Models as Modeler
Diffusion models are based on a network architecture that iteratively denoises samples, starting from random noise and gradually transforming it into structured data. This process enables them to capture complex data distributions and generate high-quality data by learning to reverse the noise-corruption process, making them highly effective as modelers.
@@
ROSIE [202] introduces text-to-image diffusion models to acquire valid data for robot to learn without additional robot data. Specifically, this method takes advantage of the art text-to-image diffusion. ROSIE inpaints diverse invisible objects for manipulation, backgrounds, and distractors with the guidance from the text. In robotics manipulation tasks like tabletop pick-and-place, GenAug [163] utilizes a small image-action demonstration dataset and produces an augmented dataset of image observations to enhance the real-world demonstration data. SynthER [164] flexibly upsamples an agent’s collected experience to train RL agents offline and online, in both proprioceptive and pixel-based environments. MTDiff [162] introduces a single diffusion model to establish large-scale multi-task offline data model, which combines Transformer backbones and prompt learning for data synthesis in multitask offline scenarios.
@@ 4.2.6 GFlowNets as Modeler
Generative Flow Networks (GFlowNets) model complex data distributions by learning to sample from reward-defined distributions, making them effective for tasks like data augmentation and simulation.
In [172], GFlowNets treat sampling as a decision-making process, using Markovian trajectories to model data distributions. This unifies various generative models, enabling efficient training and inference algorithms.
[173] extends this idea with energy-based GFlowNets (EB-GFNs), combining GFlowNets with energy functions to generate samples from energy-defined distributions. This approach improves probabilistic modeling, particularly for high-dimensional discrete data.
In [203], GFlowNets are applied to text-to-image generation. The Diffusion Alignment with GFlowNet (DAG) algorithm post-trains diffusion models by generating high-reward images with high probability, addressing the challenges of alignment in generative models.
@@ 4.2.7 Autoregressive Models as Modeler
Autoregressive models, such as PixelCNN [85] and PixelRNN [183], generate data by modeling the conditional dependencies between data points, predicting each component sequentially based on previous ones. This structure allows them to capture complex patterns in data, making them highly effective ”modelers”, especially in tasks like image generation, where they predict pixel distributions to generate coherent, high-quality images.
@@ 4.3 Generative Models as Optimizer
Recent research has increasingly approached optimization problems as sampling tasks, with generative models being used to tackle these challenges. This shift allows generative models to learn data representations that can be directly used to optimize specific objective functions, making them powerful tools for solving complex optimization problems [204]. These models can capture intricate data distributions and generate solutions that are both efficient and scalable [205]. In this context, we explore how generative models, when applied as optimizers, offer new perspectives on improving optimization strategies, surpassing traditional methods in several domains.
@@ 4.3.1 EBMs as Optimizer
Energy-Based Models (EBMs) can also function as powerful optimizers by refining candidate solutions toward optimal configurations in the solution space. The key mechanism involves learning an energy function that not only evaluates the quality of solutions but also guides the optimization process. EBMs optimize the energy landscape by employing techniques like contrastive divergence [206] or stochastic gradient descent. Rather than generating random samples, the optimization process is driven by minimizing the energy function, which allows EBMs to select the most optimal solutions according to the learned energy landscape. Methods like MCMC or Langevin dynamics are often employed to iteratively refine candidate solutions, effectively navigating the solution space to converge toward the global or local minimum.
EBMs have shown their effectiveness in various optimization tasks. One approach, SO-EBM [122], leverages EBMs to directly parameterize optimization problems through differentiable energy functions. Unlike traditional methods that rely on KKT conditions to implicitly induce an optimization layer, SO-EBM explicitly defines the optimization process by using an energy-based model, allowing for more accurate capture of the optimization landscape. This approach enables more efficient and refined optimization, particularly in complex, high-dimensional tasks.
Additionally, the Pareto-compositional energy-based model (pcEBM) [123] is designed for protein sequences optimization and sampling. In this approach, pcEBM integrates multiple objectives into a single energy function through a compositional structure that balances competing performance criteria. The model learns a multi-objective energy function that simultaneously accounts for factors such as structural stability and sequence diversity, which are critical in protein design. By optimizing this compositional energy function, pcEBM efficiently navigates the solution space to identify sequences that achieve a Pareto-optimal balance. This method enhances the optimization process, making it particularly valuable for tasks requiring the simultaneous optimization of multiple criteria, such as in protein engineering and drug design. Similarly, a multistage coarse-to-fine expanding and sampling strategy (CF-EBM) [124] also utilizes sampling as a key mechanism for optimization. While pcEBM optimizes multi-objective tasks by refining protein sequences, the coarse-to-fine method enhances sample quality by progressively refining solutions. Both methods optimize the energy landscape through sampling, highlighting the versatility of EBMs in addressing a range of optimization challenges, from multi-objective tasks to high-quality generative sampling.
@@ 4.3.2 GANs as Optimizer
Generative Adversarial Networks (GANs) have emerged as powerful optimizers, addressing challenges in high-dimensional and complex optimization problems.
He et al. [136] demonstrated how GANs can generate high-quality offspring solutions in multi-objective evolutionary algorithms, improving both solution diversity and optimality. This method alleviated the curse of dimensionality, improving optimization performance with limited data. Similarly, Sim et al. [137] utilized GANs and DCGANs for topology optimization, generating valid data for structural design optimization. Their method utilized K-means clustering analysis to select optimized solutions, showing how GANs can be leveraged for structural design optimization, efficiently navigating complex design spaces and producing feasible solutions that meet specific criteria. This highlights GANs’ flexibility in optimization tasks beyond traditional settings, extending their application to engineering and design optimization.
Furthermore, Kalehbasti et al. [138] explored the combination of Conditional GANs (C-GANs) with genetic algorithms for solving high-dimensional nonlinear optimization problems. By augmenting solutions generated by a genetic algorithm with C-GANs, they enhanced the search for optimal solutions, demonstrating that the integration of GANs with evolutionary algorithms significantly accelerates the optimization process, even in the absence of vast amounts of training data. This development underscores the growing capability of GANs to serve as powerful optimizers in diverse optimization problems.
@@ 4.3.3 VAEs as Optimizer
VAEs are effective optimizers due to their ability to learn compact probabilistic representations, which facilitate efficient exploration of large and complex decision spaces. By mapping complex high-dimensional inputs to a simpler, lower-dimensional latent space, VAEs enable more efficient exploration of the decision space for optimization tasks.
For instance, CageBO [146] employs a conditional VAE to model the distribution of viable decisions in optimization tasks. By establishing a bidirectional mapping between the initial decision space and a reduced, unconstrained latent space, CageBO is able to optimize decision-making processes more efficiently. In particular, the model uses the latent representation to navigate the decision space, enabling more effective and scalable exploration of possible solutions. This approach improves optimization performance by generating high-quality candidate solutions and reducing the computational burden typically associated with exhaustive search methods.
Similarly, Conditional Variational Autoencoders (CVAEs) are also used in optimization tasks where the decision space is complex, such as routing problems. In [147], researchers employ a CVAE to learn a latent representation of the solution space, optimizing the search for feasible solutions (see Fig. 16). By conditioning the generative process on problem-specific constraints, CVAE refines the optimization process to generate high-quality solutions within the reduced latent space.
@@ 4.3.4 Normalizing Flows as Optimizer
Normalizing Flows (NF) can also serve as powerful optimizers. The core idea behind using Normalizing Flows (NF) as an optimizer is their ability to transform simple distributions (like Gaussian or uniform) into complex ones that match the desired solution distribution. This enables the generation of high-quality solutions efficiently, with fewer iterations than traditional methods. Because NF are differentiable, they can be easily integrated with gradient-based optimization, allowing for smooth, end-to-end optimization through backpropagation.
Gabrié et al. [151] proposed an adaptive MCMC method that enhances traditional MCMC sampling by incorporating nonlocal transition kernels parameterized by normalized flows. This method enhances optimization by using NF to guide the search more efficiently. By combining local MCMC updates with NF-driven transitions, it accelerates convergence to optimal solutions, improving both efficiency and performance.
@@ 4.3.5 Diffusion Models as Optimizer
Diffusion models have shown increasing potential as optimizers. The core advantage of diffusion models lies in their iterative denoising process, which allows them to refine candidate solutions over time, making them particularly effective for black-box optimization tasks where the objective function is unknown or difficult to compute directly.
In traditional optimization problems, gradient-based methods often struggle with high-dimensional spaces, especially when the objective function is highly complex or non-convex. In contrast, diffusion models approach optimization through a process of gradually refining solutions, effectively exploring the solution space and generating promising candidates.
For instance, Denoising Diffusion Optimization Models (DDOM) [165] introduce a novel inverse approach for solving offline black-box optimization problems by leveraging diffusion models. This method models complex objective functions through a reverse diffusion process, offering an efficient way to explore and optimize in high-dimensional, poorly understood spaces. By iterating through multiple diffusion steps, DDOM can refine candidate solutions to find near-optimal results even when the objective function is opaque or difficult to compute directly.
Similarly, Li et al. [166] reformulate optimization as a conditional sampling task and propose a reward-directed conditional diffusion model as Fig. 17. This model is trained using mixed data to sample near-optimal solutions conditioned on predicted rewards, guiding the optimization process in a direction that maximizes the objective function’s expected value. This reward-directed approach allows the model to focus on the most promising areas of the solution space, further enhancing its efficiency in optimization tasks.
@@
Another approach, DiffOPT [167], reframes the optimization problem by drawing samples from a joint Boltzmann distribution determined by both the objective function and the learned data distribution. This method incorporates a two-stage framework: the first stage uses a guided diffusion process to initialize the optimization, while the second stage applies Langevin dynamics to further refine the solution. The combination of these two stages allows for both efficient exploration and precise refinement, making DiffOPT particularly effective for complex optimization tasks.
@@ 4.3.6 GFlowNets as Optimizer
Generative Flow Networks (GFlowNets) have emerged as powerful optimizers, effectively navigating complex solution spaces through a unique generative flow framework. By constructing solutions step by step according to a learned flow, GFlowNets can efficiently explore and exploit the underlying structure of optimization problems. The flow-based nature of GFlowNets enables them to generate diverse candidate solutions while maintaining high sample efficiency, which is crucial in scenarios where the evaluation of solutions is computationally expensive. Their ability to learn from both on-policy and off-policy data further enhances their adaptability and performance across a wide range of optimization tasks.
[207] explores the application of GFlowNets to combinatorial optimization (CO) problems, which are often NP-hard and challenging for exact algorithms. The researchers design a MDP tailored to various combinatorial problems and train conditional GFlowNets to sample from the solution space. Extensive experiments on diverse CO tasks with both synthetic and realistic data have demonstrated that GFlowNet policies can effectively find high-quality solutions.
Building on this foundation, GFlowNets were extended to multi-objective optimization, addressing the challenge of balancing multiple conflicting objectives. MOGFNs [175], a novel method based on GFLowNets, can generate diverse Pareto-optimal solutions by learning to navigate the trade-offs between different objectives. This development significantly broadened the scope of GFlowNets’ applicability in optimization.
Additionally, Generative Flow Ant Colony Sampler (GFACS) [174] hierarchically combines amortized inference and parallel stochastic search. The solution sampling process is shown in Fig. LABEL:GFACS. This hybrid approach leverages the strengths of both methods: the efficient sampling and flow-based learning of GFlowNets, and the heuristic search capabilities of ACO. GFACS demonstrated significant improvements in solving combinatorial optimization problems by incorporating off-policy training and local search strategies.
Figure 18: Solution sampling mechanism of GFACS [166]. The GNN, trained with a GFlowNet loss, acts as an expert guide to select the next step in constructing a solution, like a tour in the TSP.

### v3

@@ 4.2 Generative Models as Modeler
Unlike methods like S4RL [180] or RAD [181], which perturb inputs to maintain consistency, generative models focus on learning the underlying patterns within data to generate new data that mirrors the original distribution [182].
As modelers, generative models analyze input data to capture its inherent structures and relationships [183]. They then use this understanding to generate new data, such as images or text, that resemble the original dataset. For example, DALL·E [1] create images with similar style and content to the training data. Beyond image and text generation, they are also useful for synthetic data generation in tasks like data augmentation, privacy protection, and simulations.
@@ 4.2.1 EBMs as Modeler
EBMs model data using an energy function that captures its underlying structure, without relying on explicit probability distributions. Instead, they describe data through an energy landscape, and once trained, they generate new data by minimizing the energy, aligning generated samples with the learned patterns of the original dataset.
Boltzmann Machines (BMs) [27] use contrastive divergence to generate samples that match the training data’s distribution. DEBMs [107] extend this by learning a distribution over state-action pairs, allowing reinforcement learning agents to generate and select optimal actions. Score-Based Generative Models (SGMs) [89] use score matching to learn gradients of the data distribution and generate samples by refining noisy data towards the target distribution.
@@ 4.2.2 GANs as Modeler
GANs function as models by learning the underlying data distributions through their unique adversarial framework. The generator captures the patterns in the data and uses this knowledge to generate synthetic samples, while the discriminator helps refine these samples by assessing their realism. Over time, this process allows the generator to better model the original data distribution, making it a powerful tool to generate new realistic data that reflect the learned structure of the dataset.
For instance, EGAN [121] models the relationship between states and actions, enabling the pre-training of agents and accelerating their learning. Similarly, State2Pixel (S2P) [122] generates pixel images from the agent’s state, bridging the gap between state space and image domains in reinforcement learning. S2P also enables virtual exploration of the latent image distribution through model-based state transitions. Thus, GANs excel as modelers by capturing complex data relationships and generating data that mirrors the learned patterns.
@@ 4.2.3 VAEs as Modeler
VAEs function as modelers by learning a probabilistic mapping from the data space to a latent space, where they capture the underlying variations and structure of the data. Through this process, VAEs represent complex data distributions in a more compact and continuous latent space.
For example, Han & Kim [132] use VAEs to model the subspaces of a dataset. The VAE encoder transforms data points into the latent space, and the decoder generates new data by sampling from this learned latent space and mapping it back to the original data space. This capability to model and generate data based on learned distributions allows VAEs to capture the essential features of the data while producing new, similar instances.
@@ 4.2.4 Normalizing Flows as Modeler
Normalizing Flows (NFs) are powerful modelers for complex, high-dimensional data distributions, using invertible transformations to capture intricate data structures. This flexibility makes them well-suited for tasks such as density estimation, variational inference, and posterior distribution modeling and so on.
For instance, NICE [80] uses a series of invertible transformations, specifically coupling layers, to model high-dimensional data distributions. This approach allows it to transform data into a simpler latent space, where the distribution is easier to model, while maintaining computational efficiency by ensuring the Jacobian determinant is simple to calculate. Rezende and Mohamed [73] extends NF’s use in variational inference to model posterior distributions, enabling efficient Bayesian inference. Müller et al. [184] introduces piecewise-polynomial coupling transformations, improving NF’s capacity to model complex distributions.
@@ 4.2.5 Diffusion Models as Modeler
Diffusion models generate high-quality data by iteratively denoising random noise into structured data, enabling them to capture complex data distributions and reverse the noise-corruption process, making them effective modelers.
@@
ROSIE [185] applies text-to-image diffusion models to generate valid data for robotic learning, inpainting objects for manipulation and backgrounds with text guidance. GenAug [150] enhances real-world demonstration data by augmenting a small image-action dataset in tabletop pick-and-place tasks. SynthER [151] upsamples an agent’s experience to train RL agents in both proprioceptive and pixel-based environments. MTDiff [149] introduces a multi-task offline data model using a single diffusion model with Transformer backbones and prompt learning to synthesize data in multitask scenarios.
@@ 4.2.6 GFlowNets as Modeler
GFlowNets model complex data distributions by sampling from reward-defined distributions, making them effective for tasks like data augmentation and simulation.
In [158], GFlowNets treat sampling as a decision-making process, using Markovian trajectories to model data distributions, unifying various generative models for efficient training and inference. In [159], energy-based GFlowNets (EB-GFNs) combine GFlowNets with energy functions to generate samples from energy-defined distributions, enhancing probabilistic modeling for high-dimensional discrete data. [186] applies GFlowNets to text-to-image generation, using the Diffusion Alignment with GFlowNet (DAG) algorithm to refine diffusion models and generate high-reward images, addressing alignment issues.
@@ 4.2.7 Autoregressive Models as Modeler
Autoregressive models (AMs) have been applied to reinforcement learning (RL) and planning, where their ability to model sequential dependencies proves particularly useful. As modelers, AMs learn to predict each future state or action based on the previous ones, capturing complex sequential patterns that are crucial for decision-making tasks.
For example, ALPINE [169] extends this capability by enabling agents to learn sequential policies, allowing long-term decision making by predicting future actions from past ones, which is especially effective for complex planning tasks where the order of actions matters. Similarly, ARP [170] applies AMs in RL to enhance exploration. By generating exploration policies, ARP helps agents predict diverse action trajectories, improving their ability to explore and learn in environments with sparse rewards, thus addressing the exploration-exploitation tradeoff.
@@ 4.3 Generative Models as Optimizer
Generative models have increasingly been applied to optimization problems by treating them as sampling tasks. This approach enables models to learn data representations that directly optimize objective functions, making them powerful tools for complex optimization [187]. These models can capture complex data distributions and generate efficient, scalable solutions [188]. As optimizers, generative models offer new perspectives, often outperforming traditional methods in various domains.
@@ 4.3.1 EBMs as Optimizer
Energy-Based Models (EBMs) can function as effective optimizers by refining candidate solutions toward optimal configurations. Using techniques like contrastive divergence [189] or stochastic gradient descent, EBMs minimize the energy function, offering more flexibility than traditional methods that rely on fixed rules. Methods like MCMC or Langevin dynamics refine solutions iteratively, helping EBMs navigate the solution space efficiently.
EBMs have demonstrated their strength in various optimization tasks. For example, SO-EBM [109] directly parameterizes optimization problems with differentiable energy functions, making it more accurate and efficient than traditional methods that rely on KKT conditions. Similarly, pcEBM [110], applied in protein sequence optimization, optimizes multiple objectives by learning a multi-objective energy function, allowing it to simultaneously balance structural stability and sequence diversity.
@@ 4.3.2 GANs as Optimizer
GANs have become effective optimizers, tackling high-dimensional and complex problems. Their ability to generate diverse, high-quality solutions allows them to efficiently explore large solution spaces. GANs excel in non-convex optimization, where gradient-based methods struggle, by generating realistic solutions that meet multiple constraints.
He et al. [123] showed how GANs improve multi-objective evolutionary algorithms by generating high-quality offspring, enhancing both diversity and optimality, even with limited data. Sim et al. [124] applied GANs and DCGANs in topology optimization for structural design, using K-means clustering to select optimized solutions, demonstrating GANs’ ability to navigate complex design spaces. This highlights their flexibility in engineering and design optimization. Kalehbasti et al. [125] combined Conditional GANs (C-GANs) with genetic algorithms to solve high-dimensional optimization problems. By augmenting genetic algorithm solutions with C-GANs, they accelerated the search for optimal solutions, even without large amounts of training data, showcasing GANs’ growing role as powerful optimizers in diverse tasks.
@@ 4.3.3 VAEs as Optimizer
VAEs are effective optimizers due to their ability to learn compact probabilistic representations, which allow efficient exploration of complex decision spaces. By mapping high-dimensional inputs to a simpler latent space, VAEs enhance optimization tasks by reducing computational complexity and improving solution quality.
@@
For example, CageBO [133] uses a conditional VAE to optimize decision-making by mapping decisions to a reduced latent space, allowing for more efficient exploration. This approach improves optimization by generating high-quality solutions and reducing computational costs. Similarly, Conditional VAEs (CVAEs) are applied to complex tasks like routing problems [134], where they refine the search for feasible solutions by conditioning on problem-specific constraints, improving solution quality in a reduced space (see Fig. 12).
@@ 4.3.4 Normalizing Flows as Optimizer
NFs are powerful optimizers due to their ability to transform simple distributions into complex ones, matching the desired solution distribution. This transformation allows NFs to generate high-quality solutions efficiently with fewer iterations compared to traditional methods. Being differentiable, NF can integrate seamlessly with gradient-based optimization for smooth, end-to-end optimization.
For example, Gabrié et al. [138] introduced an adaptive MCMC method that enhances traditional MCMC sampling by incorporating NF-driven non-local transitions. This approach accelerates optimization by efficiently guiding the search for optimal solutions, improving both performance and convergence speed.
@@ 4.3.5 Diffusion Models as Optimizer
Diffusion models are emerging as effective optimizers, particularly suited for black-box optimization tasks where the objective function is complex or difficult to compute. Their iterative denoising process gradually refines solutions, enabling efficient exploration of solution spaces.
In traditional optimization, gradient-based methods often struggle in high-dimensional and non-convex spaces. In contrast, diffusion models excel by gradually refining solutions, effectively navigating complex problem spaces.
For example, Denoising Diffusion Optimization Models (DDOM) [152] tackle offline black-box optimization by using reverse diffusion to model complex objective functions, refining solutions through multiple steps. Similarly, Li et al. [153] introduced a reward-directed conditional diffusion model that focuses on promising solution areas by conditioning on predicted rewards, further improving optimization efficiency.
@@
Another approach, DiffOPT [154], combines a guided diffusion process with Langevin dynamics to optimize solutions, ensuring both efficient exploration and precise refinement for complex tasks.
@@ 4.3.6 GFlowNets as Optimizer
GFlowNets have emerged as powerful optimizers that navigate complex solution spaces by generating solutions step-by-step through a learned flow. Their flow-based structure allows for diverse solution generation while maintaining high sample efficiency, crucial for expensive evaluations. GFlowNets perform well with both on-policy and off-policy data, boosting their effectiveness in tasks like combinatorial and multi-objective optimization.
For example, GFlowNets have been applied to combinatorial optimization (CO) problems, often NP-hard, where they effectively find high-quality solutions by sampling from the solution space [190]. Additionally, the extension to multi-objective optimization with MOGFNs [161] enables the generation of diverse Pareto-optimal solutions, handling trade-offs between conflicting objectives.
A hybrid approach, Generative Flow Ant Colony Sampler (GFACS) [160], combines GFlowNets with Ant Colony Optimization (ACO), further improving efficiency in combinatorial optimization by incorporating off-policy training and local search strategies.